# Learning an Adaptive and View-Invariant Vision Transformer for Real-Time UAV Tracking

### v2

@@ Abstract
Visual tracking has made significant strides due to the adoption of transformer-based models. Most state-of-the-art trackers struggle to meet real-time processing demands on mobile platforms with constrained computing resources, particularly for real-time unmanned aerial vehicle (UAV) tracking. To achieve a better balance between performance and efficiency, we introduce AVTrack, an adaptive computation framework designed to selectively activate transformer blocks for real-time UAV tracking. The proposed Activation Module (AM) dynamically optimizes the ViT architecture by selectively engaging relevant components, thereby enhancing inference efficiency without significant compromise to tracking performance. Furthermore, to tackle the challenges posed by extreme changes in viewing angles often encountered in UAV tracking, the proposed method enhances ViTs’ effectiveness by learning view-invariant representations through mutual information (MI) maximization. Two effective design principles are proposed in the AVTrack. Building on it, we propose an improved tracker, dubbed AVTrack-MD, which introduces the novel MI maximization-based multi-teacher knowledge distillation (MD) framework. It harnesses the benefits of multiple teachers, specifically the off-the-shelf tracking models from the AVTrack, by integrating and refining their outputs, thereby guiding the learning process of the compact student network. Specifically, we maximize the MI between the softened feature representations from the multi-teacher models and the student model, leading to improved generalization and performance of the student model, particularly in noisy conditions. This improvement provides a better trade-off between performance and efficiency. Extensive experiments on multiple UAV tracking benchmarks demonstrate that AVTrack-MD not only achieves performance comparable to the AVTrack but also reduces model complexity, resulting in a significant 17% increase in average tracking speed. Codes is be available at https://github.com/wuyou3474/AVTrack.
@@
Figure 1: Compared to SOTA lightweight trackers, our AVTrack-DeiT sets a new record with 84.1% precision and a speed of 256.8 FPS. Our AVTrack-MD-DeiT strikes a better trade-off with 83.6% precision and a speed of about 310.6 FPS.
@@ I Introduction
Unmanned aerial vehicle (UAV) tracking has become increasingly critical due to its diverse applications, including path planning [1], public safety [2], visual surveillance [3], environmental monitoring [4], industrial inspections [5], and agriculture [6], etc. Specifically, UAV tracking involves determining and predicting the location of a specific object in successive aerial images, which are captured by mobile cameras at relatively high altitudes. As a result, common challenging scenarios have emerged in the UAV tracking community, including fast object or UAV motion, extreme viewing angles, motion blur, low resolution, and significant occlusions where objects are obscured by other elements. Moreover, the efficiency of tracking algorithms is crucial, as mobile platforms like UAVs have limited computational resources [7, 8, 9, 10, 11]. In order to meet the unique demands of UAV applications, a practical tracker should ensure accurate tracking while running with relatively low computational and power consumption.
@@
Neither discriminative correlation filter (DCF)-based trackers nor deep learning (DL)-based trackers can achieve a satisfactory balance between performance and efficiency. DCF-based trackers offer higher efficiency but struggle with tracking accuracy and robustness, making them unsuitable for complex UAV tracking scenarios, while DL-based trackers are hindered by high computational costs and sluggish performance, limiting their practical applicability. Single-stream architecture has recently emerged as a popular strategy in DL-based trackers, seamlessly integrating feature extraction and fusion with pre-trained Vision Transformer (ViT) backbone networks. A number of representative methods, including OSTrack [12], SimTrack [13], Mixformer [14], DropMAE [15], ZoomTrack [16], AQATrack [17], EVPTrack [18], and HIPTrack [19], demonstrate the immense potential of this paradigm shift in tracking task. Motivated by this, Aba-VTrack [7] proposes an effective and efficient DL-based tracker for real-time UAV tracking based on this framework, utilizing an adaptive and background-aware token calculation technique to minimize inference time. However, the presence of unstructured access operations with a variable number of tokens results in significant time costs.
@@
In this work, we alleviate the above problems by introducing an adaptive computing framework specifically designed to selectively activate transformer blocks for real-time UAV tracking, based on the single-stream architecture driven by pre-trained transformer backbone networks. Our goal concentrates on enhancing the efficiency of ViTs by using more structured methods. Specifically, we introduce an Activation Module (AM) into each transformer block. As shown in Fig. 2 (right), to improve efficiency by reducing computation, AM utilizes only a slice of tokens from both the target template and the search image as input, it produces an activation probability that decides whether a transformer block should be activated. By adaptively trimming the ViT at the block level, the proposed method avoids unnecessary unstructured access operations, thereby reducing time consumption. The rationale behind this is based on the understanding that semantic features or relationships do not uniformly affect the tracking task across different levels of abstraction. Instead, in practice, their impact is closely tied to the characteristics of the target and the scene in which it appears. In simple scenes, such as when a target moves against a monochromatic background, efficient tracking can be achieved by utilizing the color contrast between the object and the background. In such cases, this straightforward feature is sufficient. However, in real-world scenarios, there are often numerous distractors, such as background clutter, occlusion, similar objects, and changes in viewpoint. In these challenging scenarios, trackers need to capture and analyze sufficient semantic features and relationships to successfully track the specific object. These facts demonstrate the dynamic nature of tracking needs, which are inextricably linked to the individual features of the scene and the target being tracked. The proposed AM is implemented using just a linear layer followed by a nonlinear activation function, making it a straightforward and effective module. By customizing the architecture of ViTs to suit the specific requirements of tracking tasks, our approach has the potential to achieve a satisfactory balance between performance and efficiency for UAV tracking.
@@
On the other hand, we introduce a novel approach to learning view-invariant feature representations by maximizing the mutual information between the backbone features extracted from two different views of the target, enhancing the tracker’s robustness against viewpoint changes. Specifically, mutual information is a measure that quantifies the dependence or relationship between two variables [20]. Mutual information maximization involves enhancing the mutual information between different components or variables within a system and is widely used in various computer vision tasks [21, 22, 23]. However, to the best of our knowledge, the effectiveness of this strategy in UAV tracking has not been extensively explored. In our work, to obtain view-invariant representations, the tracker learns to preserve essential information about the target regardless of viewpoint changes by maximizing the mutual information between two different views of the target. By employing this method, we call the resulting representations view-invariant representations. In theory, models trained with these representations are better at generalizing across diverse viewing conditions, making them more robust in real-world scenarios with frequent viewpoint changes. View-invariant representations are useful to trackers for effective tracking since extreme changes in viewing angles are a common challenge in UAV tracking.
@@
To enhance the single-stream paradigm for real-time UAV tracking, we conduct extensive experiments, highlight the two key design principles mentioned above, and introduce an improved version called AVTrack-MD. Specifically, we introduce a novel multi-teacher knowledge distillation (MD) framework based on MI maximization in the AVTrack. In the proposed MD, three off-the-shelf tracking models from the AVTrack serve as teachers, distilling knowledge into a more lightweight student model. The student model maintains the same structure as the teachers but features a smaller ViT backbone, with half the number of ViT blocks. In practice, multiple teacher models offer comprehensive guidance by providing diverse knowledge, which is advantageous for training the student [24, 25]. Multi-teacher knowledge distillation is an effective strategy that is widely used across various computer vision tasks, including image super-resolution [26], image classification [27], and visual retrieval [28]. However, to the best of our knowledge, its effectiveness in UAV tracking has not been extensively investigated. In our work, given the mean squared error (MSE) is sensitive to noise and outliers [29], we maximize the MI between the aggregated softened features of the multi-teacher models and the student model’s softened features, leading to improved generalization and performance of the student model, particularly in noisy conditions. Extensive experimental results demonstrate that the proposed method can produce a high-performing student network, with overall tracking performance that is comparable to or even exceeds that of teacher networks, while using significantly fewer parameters and achieving faster speeds (See Fig. 1 and Table I).
@@
This work extends our previous conference paper, i.e., AVTrack [30], accepted at ICML 2024. In this extended paper, we propose AVTrack-MD, an improved tracker that significantly enhances the efficiency of AVTrack while achieving the overall tracking performance comparable to or even superior to it. We also provide extensive experiments and detailed implementations. The new contributions of this extended work are summarized as follows:
@@
    We propose an improved version tracker, dubbed AVTrack-MD, with more lightweight architectural adaptations for UAV tracking. It significantly enhances the efficiency of AVTrack while achieving comparable or even superior performance.
@@
    We conducted experiments on six UAV tracking benchmarks, providing more comparisons with recent state-of-the-art trackers, along with a comprehensive evaluation and analysis.
@@
    We have presented more in-depth design principles, discussions, and details.
@@ II-A UAV Tracking
There are two main types of modern UAV trackers: DCF-based and CNN-based trackers [31, 32]. DCF-based trackers are favored for UAV tracking due to their remarkable efficiency on CPU, but they face difficulties in maintaining robustness under complex conditions [31, 11, 33]. Current CNN-based trackers, like those in [34, 35], show advancements in tracking precision and robustness for UAV tracking, but their efficiency lags behind DCF-based trackers considerably. Although model compression techniques, as seen in [36, 37, 38], were utilized to enhance efficiency, these trackers still face challenges associated with unsatisfactory tracking precision. A recent trend in the visual tracking community shows an increasing preference for single-stream architectures that seamlessly integrate feature extraction and correlation using pre-trained ViT backbone networks [39, 14, 12, 40, 17], demonstrating significant potential. Several representative methods, including OSTrack [12], Mixformer [14], DropMAE [15], ZoomTrack [16], AQATrack [17], EVPTrack [18], etc., demonstrate the significant success of applying this paradigm to tracking tasks. Although these frameworks are efficient due to their compact nature, very few are based on lightweight ViTs, making them impractical for real-time UAV tracking. To address this, several studies have started exploring the use of efficient ViTs built on one-stream framework for real-time UAV tracking [7, 41, 42, 43]. Specifically, ETDMT [43] builds on lightweight ViT to introduce a tracker that combines template distinction with temporal context, while Aba-ViTrack [7] enhances efficiency in real-time UAV tracking using lightweight ViTs and an adaptive background-aware token computation method. However, the variable token number in [7] necessitates unstructured access operations, leading to notable time costs. Recent research has focused on improving the efficiency of ViTs by balancing their representation capabilities with computational efficiency. Methods like lightweight ViTs, model compression, and hybrid designs [44, 45, 46, 47, 48] have been explored, but they often sacrifice accuracy or require time-consuming fine-tuning. Recent developments in efficient ViTs with conditional computation focus on adaptive inference, dynamically adjusting computational load based on input complexity to accelerate model performance. For example, DynamicViT [49] introduces control gates to selectively process tokens, while A-ViT [50] employs an Adaptive Computation Time strategy to avoid auxiliary halting networks, achieving gains in efficiency, accuracy, and token prioritization. Aba-ViTrack [7] effectively utilizes the latter, showcasing significant potential for real-time UAV tracking. In our work, we focus on improving the efficiency of ViTs for UAV tracking through more structured methods, specifically the adaptive activation of Transformer blocks for feature representation.
@@ II-B View-Invariant Feature Representation
View-invariant feature representation has garnered significant attention in the field of computer vision and image processing. This technique aims to extract features from images or visual data that remain consistent across various viewpoints or orientations, providing robustness to changes in the camera angle or scene configuration [51, 52, 53, 54]. Traditional methods typically rely on handcrafted features and geometric transformations to achieve view-invariant representations [55, 56, 54]. While effective, these traditional methods are often limited to specific scenarios with relatively simple and fixed backgrounds, making them unsuited for handling the complexity and variability of real-world visual data. Recently, many researchers have been eager to apply Convolutional Neural Networks (CNNs) and other deep architectures for extracting view-invariant representations, thanks to deep learning’s remarkable ability to extract discriminative features [57, 58, 59, 60, 61]. For instance, [57] propose the Dual-Attention Network (DANet) for view-invariant action recognition, incorporating relation-aware spatiotemporal self-attention and cross-attention modules to learn representative and discriminative action features. [62] propose a View Transformation Network (VTN) that realizes the view normalization by transforming arbitrary-view action samples to a base view to seek a view-invariant representation. These approaches leverage the capacity of deep models to capture complex patterns and variations in visual data. By ensuring that learned features are resilient to changes in viewpoint, these methods contribute to the robustness and generalization of vision-based systems. View-invariant feature representation is helpful for a variety of vision tasks, such as action recognition [63, 64, 51], pose estimation [52], human re-identification [65, 66], and object detection [67]. However, to the best of our knowledge, the integration of view-invariant representations into visual tracking frameworks remains underexplored. In our work, we make the first attempt to learn view-invariant feature representations through mutual information maximization based on ViTs, specifically tailored for UAV tracking. This marks the first instance where ViTs are employed to acquire view-invariant feature representations in the context of UAV tracking.

### v3

@@ Abstract
Transformer-based models have improved visual tracking, but most still cannot run in real time on resource-limited devices, especially for unmanned aerial vehicle (UAV) tracking. To achieve a better balance between performance and efficiency, we propose AVTrack, an adaptive computation tracking framework that adaptively activates transformer blocks through an Activation Module (AM), which dynamically optimizes the ViT architecture by selectively engaging relevant components. To address extreme viewpoint variations, we propose to learn view-invariant representations via mutual information (MI) maximization. In addition, we propose AVTrack-MD, an enhanced tracker incorporating a novel MI maximization-based multi-teacher knowledge distillation framework. Leveraging multiple off-the-shelf AVTrack models as teachers, we maximize the MI between their aggregated softened features and the corresponding softened feature of the student model, improving the generalization and performance of the student, especially under noisy conditions. Extensive experiments show that AVTrack-MD achieves performance comparable to AVTrack’s performance while reducing model complexity and boosting average tracking speed by over 17%. Codes is available at https://github.com/wuyou3474/AVTrack.
@@
Figure 1: Compared to SOTA lightweight trackers, our AVTrack-DeiT sets a new record with 84.1% precision and a speed of 256.8 FPS. Our AVTrack-MD-DeiT strikes a better trade-off with 83.6% precision and a speed of about 310.6 FPS. Note that the evaluation was conducted on a PC equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU.
@@ I Introduction
Unmanned aerial vehicle (UAV) tracking has become increasingly critical due to its diverse applications, including path planning, public safety, visual surveillance, environmental monitoring, and agriculture, etc. Specifically, UAV tracking involves determining and predicting the location of a specific object in successive aerial images, which are captured by mobile cameras at relatively high altitudes. As a result, common challenging scenarios have emerged in the UAV tracking community, including fast object or UAV motion, extreme viewing angles, motion blur, low resolution, and significant occlusions where objects are obscured by other elements. Moreover, the efficiency of tracking algorithms is crucial, as mobile platforms like UAVs have limited computational resources [1, 2, 3, 4]. In order to meet the unique demands of UAV applications, a practical tracker should ensure accurate tracking while running with relatively low computational and power consumption.
@@
Neither discriminative correlation filter (DCF)-based trackers [5, 6, 7, 2, 8, 9, 10] nor deep learning (DL)-based trackers [11, 12, 13, 3, 14, 15, 16, 17, 18, 19, 20] can achieve a satisfactory balance between performance and efficiency. DCF-based trackers offer higher efficiency but struggle with tracking accuracy and robustness, making them unsuitable for complex UAV tracking scenarios, while DL-based trackers are hindered by high computational costs and sluggish performance, limiting their practical applicability. Single-stream architecture has recently emerged as a popular strategy in DL-based trackers, seamlessly integrating feature extraction and fusion with pre-trained Vision Transformer (ViT) backbone networks. A number of representative methods [21, 22, 23, 24, 25, 26, 27, 28, 29, 30], including OSTrack [21], SimTrack [22], Mixformer [23], DropMAE [24], ZoomTrack [25], AQATrack [26], EVPTrack [27], and HIPTrack [28], demonstrate the immense potential of this paradigm shift in tracking task. Motivated by this, Aba-VTrack [1] proposes an effective and efficient DL-based tracker for real-time UAV tracking based on this framework, utilizing an adaptive and background-aware token calculation technique to minimize inference time. However, the presence of unstructured access operations with a variable number of tokens results in significant time costs.
@@
In this work, we alleviate the above problems by introducing AVTrack [31], an adaptive computation framework for real-time UAV tracking that enhances the efficiency of Transformer-based models by integrating an Activation Module (AM) into each block. This module dynamically selects relevant components during the tracking, significantly improving speed while maintaining competitive tracking performance. As shown in Fig. 2 (right), to improve efficiency by reducing computation, AM utilizes only a slice of tokens from both the target template and the search image as input, it produces an activation probability that decides whether a transformer block should be activated. By adaptively trimming the ViT at the block level, the proposed method avoids unnecessary unstructured access operations, thereby reducing time consumption. The rationale behind this is based on the understanding that semantic features or relationships do not uniformly affect the tracking task across different levels of abstraction. Instead, in practice, their impact is closely tied to the characteristics of the target and the scene in which it appears. In simple scenes, such as when a target moves against a monochromatic background, efficient tracking can be achieved by utilizing the color contrast between the object and the background. In such cases, this straightforward feature is sufficient. However, in real-world scenarios, there are often numerous distractors, such as background clutter, occlusion, similar objects, and changes in viewpoint. In these challenging scenarios, trackers need to capture and analyze sufficient semantic features and relationships to successfully track the specific object. These facts demonstrate the dynamic nature of tracking needs, which are inextricably linked to the individual features of the scene and the target being tracked. The proposed AM is implemented using just a linear layer followed by a nonlinear activation function, making it a straightforward and effective module. By customizing the architecture of ViTs to suit the specific requirements of tracking tasks, our approach has the potential to achieve a satisfactory balance between performance and efficiency for UAV tracking.
@@
Additionally, AVTrack [31] introduces a novel approach for learning view-invariant feature representations by maximizing mutual information between backbone features from two different views of the target, thereby enhancing the tracker’s robustness against viewpoint variations. Specifically, mutual information is a measure that quantifies the dependence or relationship between two variables [32]. Mutual information maximization involves enhancing the mutual information between different components or variables within a system and is widely used in various computer vision tasks [33, 34, 35]. However, to the best of our knowledge, the effectiveness of this strategy in UAV tracking has not been extensively explored. In our work, to obtain view-invariant representations, the tracker learns to preserve essential information about the target regardless of viewpoint changes by maximizing the mutual information between two different views of the target. By employing this method, we call the resulting representations view-invariant representations. In theory, models trained with these representations are better at generalizing across diverse viewing conditions, making them more robust in real-world scenarios with frequent viewpoint changes. View-invariant representations are useful to trackers for effective tracking since extreme changes in viewing angles are a common challenge in UAV tracking.
@@
To strengthen the single-stream paradigm for real-time UAV tracking, we conduct extensive experiments, highlight the two core design principles of AVTrack , and introduce an improved version called AVTrack-MD. Specifically, we introduce a novel multi-teacher knowledge distillation (MD) framework based on MI maximization into the AVTrack. In the proposed MD, three off-the-shelf tracking models from the AVTrack serve as teachers, distilling knowledge into a more lightweight student model. The student model adopts the same structure as the teacher models but uses a smaller ViT backbone, containing only half the number of blocks. In practice, multiple teacher models offer comprehensive guidance by providing diverse knowledge, which is advantageous for training the student [36, 37]. Multi-teacher knowledge distillation is an effective strategy that is widely used across various computer vision tasks, including image super-resolution [38], image classification [39], and visual retrieval [40]. However, to the best of our knowledge, its effectiveness in UAV tracking has not been extensively investigated. In our work, given that the mean squared error (MSE) is sensitive to noise and outliers [41], we maximize the MI between the aggregated softened features of the multi-teacher models and the student model’s softened features, leading to improved generalization and performance of the student model, particularly in noisy conditions. Extensive experimental results demonstrate that the proposed method can produce a high performing student network, with overall tracking performance that is comparable to or even exceeds that of teacher networks, while using significantly fewer parameters and achieving faster speeds (see Fig. 1 and Table I).
@@
This work is an extension of our previous conference paper, AVTrack [31], accepted at ICML 2024. In this extended paper, we propose AVTrack-MD, an improved tracker that significantly enhances the efficiency of AVTrack while achieving the overall tracking performance comparable to or even superior to it. We also provide extensive experiments and detailed implementations. The main contributions of our work are summarized as follows:
@@
    We propose AVTrack, an efficient real-time tracking framework that dynamically activates transformer blocks and learns view-invariant representations by maximizing mutual information between different target views, achieving promising performance while maintaining extremely fast tracking speeds.
@@
    We further propose AVTrack-MD, a lightweight variant of AVTrack designed specifically for UAV tracking. It incorporates simplified architectural designs that significantly improve computational efficiency while maintaining or even surpassing the accuracy of the original model.
@@
    We have proposed additional model variants and conducted experiments on more UAV tracking benchmarks, such as WebUAV-3M [42], UAVTrack112L [43], and UAV20L [44] , offering extensive comparisons with recent state-of-the-art trackers along with comprehensive evaluations and analyses.
  4. 4.
    We also provide more detailed discussions on design principles and implementation strategies to clearly illustrate our approach and serve as a reference for designing more effective and efficient trackers.
@@ II-A UAV Tracking
There are two main types of modern UAV trackers: DCF-based [8, 6, 45, 2, 7, 9, 10] and DL-based trackers [46, 15, 16, 3, 14]. For instance, TB-BiCF [8] introduces a bidirectional incongruity-aware correlation filter that leverages a temporary block to capture and store inter-frame information for short durations. MT-Track [16] presents an optimized multi-step temporal modeling framework that effectively utilizes temporal context from past frames to enhance UAV tracking. DCF-based trackers offer efficient performance but lack robustness due to limited feature representation, while DL-based trackers are constrained by high computational costs and sluggish performance, limiting their practical applicability. Although model compression techniques, as seen in [47], were utilized to enhance efficiency, these trackers still face challenges associated with unsatisfactory tracking precision. A recent trend in the visual tracking community shows an increasing preference for single-stream architectures that seamlessly integrate feature extraction and correlation using pre-trained ViT backbone networks [48, 23, 21, 49, 26, 30, 29], demonstrating significant potential. Several representative methods, including OSTrack [21], Mixformer [23], DropMAE [24], ZoomTrack [25], AQATrack [26], EVPTrack [27], etc., demonstrate the significant success of applying this paradigm to tracking tasks. Although these frameworks are efficient due to their compact nature, very few are based on lightweight ViTs, making them impractical for real-time UAV tracking. To address this, several studies have started exploring the use of efficient ViTs built on one-stream framework for real-time UAV tracking [1, 4, 50, 51]. Specifically, ETDMT [51] builds on lightweight ViT to introduce a tracker that combines template distinction with temporal context, while Aba-ViTrack [1] enhances efficiency in real-time UAV tracking using lightweight ViTs and an adaptive background-aware token computation method. However, the variable token number in [1] necessitates unstructured access operations, leading to notable time costs. Recent research has focused on improving the efficiency of ViTs by balancing their representation capabilities with computational efficiency. Methods like lightweight ViTs, model compression, and hybrid designs [52, 53, 54, 55] have been explored, but they often sacrifice accuracy or require time-consuming fine-tuning. Recent developments in efficient ViTs with conditional computation focus on adaptive inference, dynamically adjusting computational load based on input complexity to accelerate model performance. For example, DynamicViT [56] introduces control gates to selectively process tokens, while A-ViT [57] employs an Adaptive Computation Time strategy to avoid auxiliary halting networks, achieving gains in efficiency, accuracy, and token prioritization. Aba-ViTrack [1] effectively utilizes the latter, showcasing significant potential for real-time UAV tracking. In our work, we focus on improving the efficiency of ViTs for UAV tracking through more structured methods, specifically the adaptive activation of transformer blocks for feature representation.
@@ II-B View-Invariant Feature Representation
View-invariant feature representation has garnered significant attention in the field of computer vision and image processing. This technique aims to extract features from images or visual data that remain consistent across various viewpoints or orientations, providing robustness to changes in the camera angle or scene configuration [58, 59, 60, 61]. Traditional methods typically rely on handcrafted features and geometric transformations to achieve view-invariant representations [62, 63, 61]. While effective, these traditional methods are often limited to specific scenarios with relatively simple and fixed backgrounds, making them unsuited for handling the complexity and variability of real-world visual data. In recent years, Convolutional Neural Networks (CNNs) and other deep learning architectures have been widely adopted to extract view-invariant representations, due to their proven ability to learn highly discriminative features from complex data [58, 64, 65, 66]. For instance, Kumie et al. [58] proposed the Dual-Attention Network (DANet) for view-invariant action recognition, which integrates relation-aware spatiotemporal self-attention and cross-attention modules to effectively learn representative and discriminative action features. Gao et al. [65] introduced the View Transformation Network (VTN) that realizes the view normalization by transforming arbitrary-view action samples to a base view to seek a view-invariant representation. These approaches leverage the capacity of deep models to capture complex patterns and variations in visual data. By ensuring that learned features are resilient to changes in viewpoint, these methods contribute to the robustness and generalization of vision-based systems. View-invariant feature representation is helpful for a variety of vision tasks, such as action recognition [62, 58], pose estimation [59], human re-identification [67], and object detection [68]. However, to the best of our knowledge, the integration of view-invariant representations into visual tracking frameworks remains underexplored. In our work, we make the first attempt to learn view-invariant feature representations through mutual information maximization based on ViTs, specifically tailored for UAV tracking. This marks the first instance where ViTs are employed to acquire view-invariant feature representations in the context of UAV tracking.